OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (4, - 6) m = 3/5

OpenStudy (anonymous):

use the point slope formula.

OpenStudy (katrinakaif):

Use the equation (y-y1) = m (x-x1) as (4, -6) = (x1,y1)

OpenStudy (anonymous):

If you could help me with one other thing here it is : Find the domain and range of the relation. Age of Person : 64, 36, 29, 29 Books read: 42,37,37,17

OpenStudy (anonymous):

\[y-(-6)=\frac{3}{5}(x-4)\]\[y+6=\frac{3}{5}x-\frac{12}{5}\]\[y=\frac{3}{5}x-\frac{12}{5}-6\]\[y=\frac{3}{5}x-\frac{12-30}{5}\]\[y=\frac{3}{5}x-\frac{42}{5}\]